OpenStudy (anonymous):

Give the equation of a line in slope-intercept from that passes through the point (6,-9) and is parallel to the line y= 2/3x+7

OpenStudy (katrinakaif):

When an equation needs to be parallel - THEY HAVE THE SAME SLOPE In this Equation, y = 2/3x + 7 (y = mx + b) m or 2/3 is the slope. Next Since you know the slope, use teh equation (y-y1) =m(x-x1) as points (6, -9) =(x1,y1) M = 2/3 (slope) and SOLVE =)

OpenStudy (anonymous):

if it is paraller then the slope must be the same ( 2/3) and should passes trhough the point (6, -9) so the only thing you need to find is the b ( intersect with y-axis) y = mx +b -9 = (2/3) * 6 +b b = -13 so the answer is: y=(2/3)x -13